# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

### v1

@@ 2.1. Correlation Functions and a priori estimates
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\mathbb{E}\big{[}\|\theta_{n}(r)\|_{2}^{q}\big{]}^{2/q}\,\mathrm{d}r\bigg{)}^{q/2}$
@@
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{dq/2}\mathbb{E}\bigg{(}\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big{\|}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r))\big{\|}_{H^{-\alpha-1+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim n^{dq/2}\mathbb{E}\bigg{(}\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}.$
@@
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta}}^{2}$ $\lesssim\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta)}\Big{(}\big{|}\widehat{\theta_{n}}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\,\mathrm{d}\xi$
$\lesssim_{\alpha,\delta}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2};$
@@
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}\|\varphi\|_{2}^{q}\bigg{(}\int_{s}^{t}\frac{1}{|t-r|^{1-\delta}}\mathbb{E}\big{[}\|\theta_{n}(r)\|_{2}^{q}\big{]}^{2/q}\,\mathrm{d}r\bigg{)}^{q/2}$
@@
thus for any $\varepsilon>0$, we deduce from [16, Corollary A.5] that the laws of $\{\mathcal{X}_{n}\}_{n}$ are tight in $C^{\gamma}([0,T];H^{-\alpha-\varepsilon}_{\rm loc})$. By the arbitrariness of $\alpha>d/2$ and $\varepsilon>0$, we conclude the desired assertion. ∎
@@
$\mathrm{d}\mathcal{X}_{n}(t)=(\kappa+\nu)\Delta\mathcal{X}_{n}(t)\,\mathrm{d}t-n^{d/2}\nabla\cdot(V^{n}(\,\mathrm{d}t)\theta_{n}(t));$
@@
$M_{n}(t)=n^{d/2}\int_{0}^{t}\nabla\cdot(\theta_{n}(s)V^{n}(\,\mathrm{d}s))=n^{d/2}\sum_{k}\int_{0}^{t}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(s))\,\mathrm{d}B^{n}_{k}(s).$
@@
We consider the quantity $\Xi_{n}=\big{(}\mathcal{X}_{n},M_{n},\theta_{n},\{B^{n}_{k}\}_{k}\big{)}$ which takes values in the product space
$\mathcal{S}=C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})\times C^{\gamma}([0,T];H^{-\beta}_{\rm loc})\times C([0,T],H^{-\alpha})\times C([0,T],\mathbb{R}^{\mathbb{N}}).$
@@
    $\tilde{\Xi}=\big{(}\tilde{\mathcal{X}},\tilde{M},\tilde{\theta},\{\tilde{B}^{k}\}_{k}\big{)}\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu$ and $\tilde{\Xi}_{n}=\big{(}\tilde{\mathcal{X}}_{n},\tilde{M}_{n},\tilde{\theta}_{n},\{\tilde{B}^{n}_{k}\}_{k}\big{)}\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu_{n}$ for any $n\geq 1$;
@@
Item (a) implies that $\tilde{\theta}=\bar{\theta}$ solves the deterministic heat equation (1.13); moreover, for any $n\geq 1$, the following equation holds in the distributional sense:
@@
$=(\kappa+\nu)\Delta\tilde{\mathcal{X}}_{n}(t)\,\mathrm{d}t-n^{d/2}\sum_{k}\nabla\cdot(\sigma^{n}_{k}\tilde{\theta}_{n}(s))\,\mathrm{d}\tilde{B}^{n}_{k}(t).$
@@
The process $\{\tilde{M}_{t}\}_{t\in[0,T]}$ is an $H^{-\beta}$-valued martingale w.r.t. the filtration $\tilde{\mathcal{F}}_{t}=\sigma\big{(}\tilde{\theta}(s),\{\tilde{B}_{k}(s)\}_{k}:s\leqslant t\big{)}$.
@@
Let $\tilde{\mathcal{F}}^{n}_{t}=\sigma\big{(}\tilde{\theta}_{n}(s),\{\tilde{B}^{n}_{k}(s)\}_{k}:s\leqslant t\big{)}$. We know that $\tilde{M}_{n}$ is an $H^{-\beta}$-valued continuous martingale w.r.t. $\{\tilde{\mathcal{F}}^{n}_{t}\}_{t}$; thus for any $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and any bounded continuous functional $G:C([0,s],H^{-\zeta})\times C([0,s],\mathbb{R}^{\mathbb{N}})\to\mathbb{R}$, we have
$\tilde{\mathbb{E}}\big{[}\big{\langle}\tilde{M}_{n}(t)-\tilde{M}_{n}(s),\phi\big{\rangle}\,G\big{(}\tilde{\theta}_{n}(\cdot),\{\tilde{B}^{n}_{k}(\cdot)\}_{k}\big{)}\big{]}=0.$
@@
By the arbitrariness of $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ and the functional $G$, we deduce that $\tilde{M}$ is an $H^{-\beta}$-valued martingale w.r.t. $\{\tilde{\mathcal{F}}_{t}\}_{t}$. ∎
@@
where we have written $\tilde{\theta}_{n}(s)$ as $\tilde{\theta}^{n}_{s}$ to save space. We have
@@
$\langle\tilde{M}_{n},\phi\rangle](t)\rightarrow\int_{[0,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x)\big{)}^{2}\nabla\phi(x)^{T}V_{\rm eff}^{2}\nabla\phi(x)\,\mathrm{d}x\,\mathrm{d}r,$
@@
where $Q^{n}(z)=Q(nz),\,z\in\mathbb{R}^{d}$. Writing $Q_{n}(z)=n^{d}Q(nz)$ yields the expression
@@
First, we show that $\mathbb{E}[|T_{1}|]=O(\varepsilon)$. Indeed we have
$\mathbb{E}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\int_{\mathbb{R}^{2d}}\mathbb{E}\big{[}|\tilde{\theta}^{n}_{s}(x)\tilde{\theta}^{n}_{s}(y)|\big{]}|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s.$
@@
$\mathbb{E}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\int_{\mathbb{R}^{2d}}q_{cs+1}(x)q_{cs+1}(y)|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s=O(\varepsilon),$
@@
Since $\varepsilon$ is arbitrarily small, we only need to handle the term $T_{2}$. Recall that we denote by $\Psi(t,x)$ the space-time stationary solution to (1.7), constructed in [10]. We observe that
$\mathbb{E}\bigg{[}\bigg{|}T_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}\tilde{\theta}^{n}_{s}(x)q_{s}\ast\varphi(y)\Psi(n^{2}s,ny)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s\bigg{|}\bigg{]}\rightarrow 0,$
@@
$\mathbb{E}\bigg{[}\bigg{|}\tilde{T}_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}q_{s}\ast\varphi(x)\Psi(n^{2}s,nx)q_{s}\ast\varphi(y)\Psi(n^{2}s,ny)$
@@
$\mathcal{T}_{2}=\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}\Psi_{s}^{n}(x)\Psi_{s}^{n}(y)q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $\Psi_{s}^{n}(x):=\Psi(n^{2}s,nx)$. To calculate the limit, we observe that
$\mathbb{E}[\mathcal{T}_{2}]=\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}w(n(x-y))q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $w$ is the spatial covariance function of $\Psi$. From [10, Corollary 3.2, Proposition 2.3] we can deduce that
$\mathbb{E}[\mathcal{T}_{2}]\rightarrow\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s,$
where $V_{\rm eff}$ is defined in (1.11). Observe that this shows that $\mathbb{E}[T_{2}]$ converges to the same limit.
@@
$\mathbb{E}\bigg{[}\bigg{|}\mathcal{T}_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg{|}^{2}\bigg{]}\rightarrow 0.$
@@
(2.22) $\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}\to\bigg{(}\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg{)}^{2}\quad\mbox{as }n\to\infty.$
@@
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $=\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}w^{(4)}_{n}(r,s,x_{1:4})q_{r}\ast\varphi(x_{1:2})q_{s}\ast\varphi(x_{3:4})$
@@
$w^{(4)}_{n}(r,s,x_{1:4})$ $=\mathbb{E}\big{[}\Psi^{n}_{r}(x_{1})\Psi^{n}_{r}(x_{2})\Psi^{n}_{s}(x_{3})\Psi^{n}_{s}(x_{4})\big{]}.$
@@
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $=\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}\tilde{w}^{(4)}_{n}(r,s,x_{1:4})q_{r}\ast\varphi(x_{2}+x_{1}/n)q_{r}\ast\varphi(x_{2})q_{s}\ast\varphi(x_{4}+x_{3}/n)q_{s}\ast\varphi(x_{4})$
@@
By Lemma 2.13, proved below, and the dominated convergence theorem, taking limit $n\to\infty$ yields
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $\to\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}w(x_{1})w(x_{3})\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\big{(}q_{s}\ast\varphi(x_{4})\big{)}^{2}$
$\quad\times\nabla\phi(x_{2})^{T}Q(x_{1})\nabla\phi(x_{2})\,\nabla\phi(x_{4})^{\ast}Q(x_{3})\nabla\phi(x_{4})\,\mathrm{d}x_{1:4}\,\mathrm{d}r\,\mathrm{d}s$
$=\bigg{[}\int_{[\varepsilon,t]}\int_{\mathbb{R}^{2d}}\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\nabla\phi(x_{2})^{T}(wQ)(x_{1})\nabla\phi(x_{2})\,\mathrm{d}x_{1:2}\,\mathrm{d}r\bigg{]}^{2}$
$=\bigg{[}\int_{[\varepsilon,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\nabla\phi(x_{2})^{T}V_{\rm eff}^{2}\nabla\phi(x_{2})\,\mathrm{d}x_{2}\,\mathrm{d}r\bigg{]}^{2},$
@@
$\mathcal{X}_{t}(\phi)=(k+\nu)\int_{0}^{t}\mathcal{X}_{s}(\Delta\phi)ds+\int_{0}^{t}\langle\nabla\phi,\bar{\theta}_{s}\,\xi(\,\mathrm{d}s)\rangle,$
@@
We conclude this section with the proof of Lemma 2.13, referenced above.

### v2

@@ 2.1. Correlation Functions and a priori estimates
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg(\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big\|\widehat{Q}\big\|_{\infty}\|\theta^{n}_{r}\|_{2}^{2}\,\mathrm{d}r\bigg)^{q/2}$
$\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\bigg(\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big\|\widehat{Q}\big\|_{\infty}\mathbb{E}\big[\|\theta^{n}_{r}\|_{2}^{q}\big]^{2/q}\,\mathrm{d}r\bigg)^{q/2}$
@@
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{dq/2}\mathbb{E}\bigg(\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big\|\nabla\cdot(\sigma^{n}_{k}\theta^{n}_{r})\big\|_{H^{-\alpha-1+\delta}}^{2}\,\mathrm{d}r\bigg)^{q/2}$
$\lesssim n^{dq/2}\mathbb{E}\bigg(\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg)^{q/2}.$
@@
$\sum_{k}\big\|\sigma^{n}_{k}\theta^{n}_{r}\big\|_{H^{-\alpha+\delta}}^{2}$ $\lesssim\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta)}\Big(\big|\widehat{\theta^{n}}(r)\big|^{2}\ast{\rm Tr}\big(\widehat{Q^{n}}\big)\Big)(\xi)\,\mathrm{d}\xi$
$\lesssim_{\alpha,\delta}n^{-d}\big\|\widehat{Q}\big\|_{\infty}\|\theta^{n}_{r}\|_{2}^{2};$
@@
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim\big\|\widehat{Q}\big\|_{\infty}^{q/2}\|\varphi\|_{2}^{q}\bigg(\int_{s}^{t}\frac{1}{|t-r|^{1-\delta}}\mathbb{E}\big[\|\theta^{n}_{r}\|_{2}^{q}\big]^{2/q}\,\mathrm{d}r\bigg)^{q/2}$
@@
thus for any $\varepsilon>0$, we deduce from [24, Corollary A.5] that the laws of $\{\mathcal{X}_{n}\}_{n}$ are tight in $C^{\gamma}([0,T];H^{-\alpha-\varepsilon}_{\rm loc})$. By the arbitrariness of $\alpha>d/2$ and $\varepsilon>0$, we conclude the desired assertion. ∎
@@
$\mathrm{d}\mathcal{X}_{n}(t)=(\kappa+\nu)\Delta\mathcal{X}_{n}(t)\,\mathrm{d}t-n^{d/2}\nabla\cdot(V^{n}(\,\mathrm{d}t)\theta^{n}_{t});$
@@
$M_{n}(t)=n^{d/2}\int_{0}^{t}\nabla\cdot(\theta^{n}(s)V^{n}(\,\mathrm{d}s))=n^{d/2}\sum_{k}\int_{0}^{t}\nabla\cdot(\sigma^{n}_{k}\theta^{n}(s))\,\mathrm{d}B^{n}_{k}(s).$
@@
We fix $\alpha>d/2$, $\gamma\in(0,1/2)$ and $\beta>1+d/2$ and we consider the quantity $\Xi_{n}=\big(\mathcal{X}_{n},M_{n},\theta^{n},\{B_{k}\}_{k},\Psi\big)$which takes values in the product space
$\mathcal{S}=C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})\times C^{\gamma}([0,T];H^{-\beta}_{\rm loc})\times C([0,T];H^{-\alpha})\times C([0,T];\mathbb{R})^{\mathbb{N}}\times L^{2}_{loc}([0,T]\times\mathbb{R}^{d}).$
@@
    $\tilde{\Xi}=\big(\tilde{\mathcal{X}},\tilde{M},\tilde{\theta},\{\tilde{B}_{k}\}_{k},\tilde{\Psi}\big)\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu$ and $\tilde{\Xi}_{n}=\big(\tilde{\mathcal{X}}_{n},\tilde{M}_{n},\tilde{\theta}^{n},\{\tilde{B}^{n}_{k}\}_{k},\tilde{\Psi}_{n}\big)\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu_{n}$ for any $n\geq 1$;
@@
[…]
Item (a) above implies that $\tilde{\theta}=\bar{\theta}$ solves the deterministic heat equation (1.13); moreover, for any $n\geq 1$, the following stochastic equation holds in the distributional sense:
@@
$=(\kappa+\nu)\Delta\tilde{\mathcal{X}}_{n}(t)\,\mathrm{d}t-n^{d/2}\sum_{k}\nabla\cdot(\sigma^{n}_{k}\tilde{\theta}^{n}(s))\,\mathrm{d}\tilde{B}^{n}_{k}(t).$
@@
The process $\{\tilde{M}_{t}\}_{t\in[0,T]}$ is an $H^{-\beta}$-valued martingale w.r.t. the filtration $\tilde{\mathcal{F}}_{t}$.
@@
We know that $\tilde{M}_{n}$ is an $H^{-\beta}$-valued continuous martingale w.r.t. $\{\tilde{\mathcal{F}}^{n}_{t}\}_{t}$; thus for any $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and any bounded continuous functional $G:C([0,s],H^{-\zeta})\times C([0,s],\mathbb{R}^{\mathbb{N}})\to\mathbb{R}$, we have
$\tilde{\mathbb{E}}\big[\big\langle\tilde{M}_{n}(t)-\tilde{M}_{n}(s),\phi\big\rangle\,G\big(\tilde{\theta}^{n}(\cdot),\{\tilde{B}^{n}_{k}(\cdot)\}_{k}\big)\big]=0.$
@@
By the arbitrariness of $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ and the functional $G$, we deduce that $\tilde{M}$ is an $H^{-\beta}$-valued martingale w.r.t. $\{\tilde{\mathcal{F}}^{0}_{t}\}_{t}$ and hence also w.r.t. $\{\tilde{\mathcal{F}}_{t}\}_{t}$. ∎
@@
where we have written $\tilde{\theta}^{n}(s)$ as $\tilde{\theta}^{n}_{s}$ to save space. We have
@@
$\langle\tilde{M}_{n},\phi\rangle](t)\rightarrow\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\big(q_{r}\ast\varphi(x)\big)^{2}\nabla\phi(x)^{\ast}V_{\rm eff}^{2}\nabla\phi(x)\,\mathrm{d}x\,\mathrm{d}r,$
@@
where the superscript $\ast$ means transposition of vectors and $Q^{n}(z)=Q(nz),\,z\in\mathbb{R}^{d}$. Writing $Q_{n}(z)=n^{d}Q(nz)=n^{d}Q^{n}(z)$ yields the expression
@@
First, we show that $\tilde{\mathbb{E}}[|T_{1}|]=O(\varepsilon)$. Indeed we have
$\tilde{\mathbb{E}}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\!\!\int_{\mathbb{R}^{2d}}\mathbb{E}\big[|\tilde{\theta}^{n}_{s}(x)\tilde{\theta}^{n}_{s}(y)|\big]|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s.$
@@
$\tilde{\mathbb{E}}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\!\!\int_{\mathbb{R}^{2d}}q_{cs+1}(x)q_{cs+1}(y)|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s=O(\varepsilon),$
@@
Since $\varepsilon$ is arbitrarily small, we only need to handle the term $T_{2}$. Recall that we denote by $\Psi(t,x)$ the space-time stationary solution to (1.7), constructed in [17]. We observe that, using the estimate in Lemma 2.4,
$\tilde{\mathbb{E}}\bigg[\bigg|T_{2}-\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{2d}}\tilde{\theta}^{n}_{s}(x)q_{s}\ast\varphi(y)\tilde{\Psi}_{n}(n^{2}s,ny)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s\bigg|\bigg]\rightarrow 0,$
@@
$\tilde{\mathbb{E}}\bigg[\bigg|\tilde{T}_{2}-\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{2d}}q_{s}\ast\varphi(x)\tilde{\Psi}_{n}(n^{2}s,nx)q_{s}\ast\varphi(y)\tilde{\Psi}_{n}(n^{2}s,ny)$
@@
$\mathcal{T}_{2}=\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{2d}}\tilde{\Psi}_{n}^{n}(s,x)\tilde{\Psi}_{n}^{n}(s,y)q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $\tilde{\Psi}_{n}^{n}(s,x):=\tilde{\Psi}_{n}(n^{2}s,nx)$. To calculate the limit, we observe that
$\tilde{\mathbb{E}}[\mathcal{T}_{2}]=\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{2d}}w_{n}(x-y)q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $w_{n}$ is the spatial covariance function of $\tilde{\Psi}_{n}^{n}$ (that is, of $\Psi(n^{2}t,nx)$). From [17, Corollary 3.2, Proposition 2.3] we can deduce that
$\tilde{\mathbb{E}}[\mathcal{T}_{2}]\rightarrow\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s,$
where $V_{\rm eff}$ is defined in (1.11). Observe that this shows that $\tilde{\mathbb{E}}[T_{2}]$ converges to the same limit.
@@
$\tilde{\mathbb{E}}\bigg[\bigg|\mathcal{T}_{2}-\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg|^{2}\bigg]\rightarrow 0.$
@@
(2.22) $\tilde{\mathbb{E}}\big[\mathcal{T}_{2}^{2}\big]\to\bigg(\int_{\varepsilon}^{t}\!\!\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg)^{2}\quad\mbox{as }n\to\infty.$
@@
$\tilde{\mathbb{E}}\big[\mathcal{T}_{2}^{2}\big]$ $=\int_{[\varepsilon,t]^{2}}\!\int_{\mathbb{R}^{4d}}w^{(4)}_{n}(r,s,x_{1:4})(q_{r}\ast\varphi)^{\otimes 2}(x_{1:2})(q_{s}\ast\varphi)^{\otimes 2}(x_{3:4})$
@@
$w^{(4)}_{n}(r,s,x_{1:4})$ $=\mathbb{E}\big[\tilde{\Psi}_{n}^{n}(r,x_{1})\tilde{\Psi}_{n}^{n}(r,x_{2})\tilde{\Psi}_{n}^{n}(s,x_{3})\tilde{\Psi}_{n}^{n}(s,x_{4})\big].$
@@
$\tilde{\mathbb{E}}\big[\mathcal{T}_{2}^{2}\big]$ $=\int_{[\varepsilon,t]^{2}}\!\int_{\mathbb{R}^{4d}}\tilde{w}^{(4)}_{n}(r,s,x_{1:4})q_{r}\ast\varphi(x_{2}+x_{1}/n)q_{r}\ast\varphi(x_{2})q_{s}\ast\varphi(x_{4}+x_{3}/n)q_{s}\ast\varphi(x_{4})$
@@
By Lemma 2.14, stated and proved below, and the dominated convergence theorem, taking limit $n\to\infty$ yields
$\tilde{\mathbb{E}}\big[\mathcal{T}_{2}^{2}\big]$ $\to\int_{[\varepsilon,t]^{2}}\!\int_{\mathbb{R}^{4d}}w(x_{1})w(x_{3})\big(q_{r}\ast\varphi(x_{2})\big)^{2}\big(q_{s}\ast\varphi(x_{4})\big)^{2}$
$\quad\times\nabla\phi(x_{2})^{\ast}Q(x_{1})\nabla\phi(x_{2})\,\nabla\phi(x_{4})^{\ast}Q(x_{3})\nabla\phi(x_{4})\,\mathrm{d}x_{1:4}\,\mathrm{d}r\,\mathrm{d}s$
$=\bigg[\int_{[\varepsilon,t]}\!\int_{\mathbb{R}^{2d}}\big(q_{r}\ast\varphi(x_{2})\big)^{2}\nabla\phi(x_{2})^{\ast}(wQ)(x_{1})\nabla\phi(x_{2})\,\mathrm{d}x_{1:2}\,\mathrm{d}r\bigg]^{2}$
$=\bigg[\int_{[\varepsilon,t]}\!\int_{\mathbb{R}^{d}}\big(q_{r}\ast\varphi(x_{2})\big)^{2}\nabla\phi(x_{2})^{\ast}V_{\rm eff}^{2}\nabla\phi(x_{2})\,\mathrm{d}x_{2}\,\mathrm{d}r\bigg]^{2},$
@@
$\mathcal{X}_{t}(\phi)=(k+\nu)\int_{0}^{t}\mathcal{X}_{s}(\Delta\phi)ds+\int_{0}^{t}\langle\nabla\phi,V_{\rm eff}\,\bar{\theta}_{s}\,\xi(\,\mathrm{d}s)\rangle,$
@@
We conclude this section with the statement and proof of Lemma 2.14, referenced above.